# ProEx: A Unified Framework Leveraging Large Language Model with Profile Extrapolation for Recommendation

### v1

@@ Abstract.
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: Proceedings of 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 9–13, 2026; Jeju, Koreaisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 2.3.1. Multi-faceted Profile Construction
It is important to note that, since we aim to characterize the user from diverse perspectives and mitigate the representational collapse caused by popularity bias, we require the LLM in $\mathcal{F}_{4}$ to generate NPs that differ entirely from OP in terms of wording and expression. Upon completing the reasoning process, we can obtain the user $u$’s profile set $\mathcal{P}_{u}=\{s_{u,1},s_{u,2},...,s_{u,\mathcal{K}}\}$, where $\mathcal{K}$ is the set of profiles. $s_{u,1}$ denotes user $u$’s OP, and $\{s_{u,2},...,s_{u,|\mathcal{K}}|\}$ correspond to the $|\mathcal{K}|-1$ NPs of user $u$. For the item side, we have a similar definition for $\mathcal{P}_{i}=\{s_{i,1},s_{i,2},...,s_{i,|\mathcal{K}|}\}$.
@@
(5) $\mathbf{C}_{u}=\{\mathbf{c}_{u,1},\mathbf{c}_{u,2},...,\mathbf{c}_{u,|\mathcal{K}|}\},\ \text{where}\ \mathbf{c}_{u,k}=f_{\text{text}}(s_{u,k})\in\mathbb{R}^{1\times d_{s}}.$
@@ 2.3.2. Profile Extrapolation
In practice, we perform a quantitative analysis with $|\mathcal{K}|=4$, where an original profile (i.e., OP) and three newly generated profiles (denoted as NP #1, NP #2, and NP #3) are obtained for each user based on multi-aspect profile construction. Fig. 4(a) illustrates the distribution of the original and newly generated profiles in the language space $\mathcal{Y}$. We observe that the original profiles (blue dots) are predominantly clustered around a central region (blue circle), reflecting a strong popularity bias. In contrast, the newly generated profiles are more widely dispersed throughout the feature space, minimizing their similarity to the original profiles. This observation demonstrates the effectiveness of the proposed CoT-based profile construction process. By occupying a broader region and expanding the feature space boundary, this process gives rise to profile extrapolation.
@@
(7) $\mathcal{L}_{\text{reg}}=\sum_{k\in\mathcal{K}}\text{log}\left(1+\text{exp}(1/\tau)\sum_{k^{\prime}\in\mathcal{K},k\neq k^{\prime}}\text{exp}\left(\tilde{\mathbf{c}}_{u,k}^{\top}\tilde{\mathbf{c}}_{u,k^{\prime}}/\tau\right)\right),$
@@
Given the difficulty of evaluating the quality of profiles, it remains uncertain whether all newly generated profiles positively contribute to the recommendation. Based on this concern, we adopt Mult-VAE (liang2018variational) as the base model and incorporate each profile individually into its training process on Amazon-Book dataset (zhang2025towards), as shown in Fig. 4(b). As is evident, there exists a substantial performance disparity among different profiles, with NP #2 even leading to a significant degradation in recommendation performance. Combined with Fig. 4(a), although we explicitly enforce reduced semantic similarity in the newly generated profiles, this process may inadvertently compromise the original semantic information. Therefore, it is necessary to reconsider how to effectively integrate these profiles $\{\tilde{\mathbf{c}}_{u,k}\}_{k=1}^{|\mathcal{K}|}$ for recommendation task.
@@
(8) $\tilde{\mathbf{c}}_{u}^{e}=\vartheta_{1}^{e}\tilde{\mathbf{c}}_{u,1}+\vartheta^{e}_{2}\tilde{\mathbf{c}}_{u,2}+...+\vartheta^{e}_{|\mathcal{K}|}\tilde{\mathbf{c}}_{u,|\mathcal{K}|};\quad\text{s.t.}\sum_{k\in\mathcal{K}}\vartheta_{k}^{e}=1,$
where LLM-generated profiles $\{\tilde{\mathbf{c}}_{u,k}\}_{k=1}^{|\mathcal{K}|}$ are weighted by $\{\vartheta^{e}_{k}\}_{k=1}^{|\mathcal{K}|}$.
For generative methods, according to Eq. 6, we model the mean $\tilde{\boldsymbol{\mu}}_{u,k}$ and variance $\tilde{\boldsymbol{\sigma}}^{2}_{u,k}$ over each user $u$’s $\mathcal{K}$ profiles. Therefore, we consider them separately:
(9) $\tilde{\boldsymbol{\mu}}_{u}^{e}=\vartheta_{1}^{e}\tilde{\boldsymbol{\mu}}_{u,1}+\vartheta^{e}_{2}\tilde{\boldsymbol{\mu}}_{u,2}+...+\vartheta^{e}_{|\mathcal{K}|}\tilde{\boldsymbol{\mu}}_{u,|\mathcal{K}|};\quad\text{s.t.}\sum_{k\in\mathcal{K}}\vartheta_{k}^{e}=1,$
@@
In each environment $e$, the construction process for discriminative and generative methods is identical. Concretely, given a fixed set of profiles, their respective contributions are controlled by a set of coefficients $\{\vartheta^{e}_{k}\}_{k=1}^{|\mathcal{K}|}$:
(10) $\{\vartheta_{1}^{e},\vartheta_{2}^{e},...,\vartheta_{|\mathcal{K}|}^{e}\}\sim\text{Dirichlet}(\alpha_{1},\alpha_{2},...,\alpha_{|\mathcal{K}|}),$
where $\{\alpha_{k}\}$ is a set of shape parameters (zhang2025mixrec) in the Dirichlet distribution. Here, we introduce a sampling-based strategy to generate entirely different combinations of coefficients $\{\vartheta^{e}_{k}\}_{k=1}^{|\mathcal{K}|}$ in each environment $e$, thereby introducing variability across environments to enhance generalization.
@@ 2.4. ProEx Framework
where $\lambda_{2}$ and $\lambda_{3}$ are also adjustable weights. For each profile $k\in\mathcal{K}$, we apply contrastive regularization $\mathcal{L}_{\text{reg}}$ (cf. Eq. 7) to further enable profile extrapolation, thereby exploring a broader preference space. Finally, we minimize the variance across all environments $\mathcal{E}$ to reduce inter-environment discrepancies, thereby extracting the invariant components of user preferences and maintaining a more stable training process (krueger2021out). The complete training process of ProEx is presented in Appendix A.1.
@@ 3.1.1. Datasets
We utilize three widely adopted recommendation datasets — Amazon-Book, Yelp, and Steam (ren2024representation). Additional details can be found in Appendix A.3.1.
@@ 3.1.4. Implementation Details
We implement ProEx by PyTorch [^2] https://github.com/BlueGhostYi/ProRec. All models are initialized using Xavier initialization (glorot2010understanding) and optimized with the Adam optimizer (kingma2014adam). The architectures of the base models adhere to the default configurations. The default batch size is 4,096 for discriminative methods, in contrast to 1,024 for generative methods. For all LLM-enhanced methods, we use OpenAI’s GPT-3.5-turbo as the base LLM and text-embedding-ada-002 (neelakantan2022text) for text embeddings to ensure a fair comparison. To evaluate the recommendation performance, we use the metrics Recall@N and NDCG@N (he2017neural). The full-ranking evaluation (he2020lightgcn) is applied for each user. Early stopping is activated when Recall@20 on the validation set does not improve for 20 consecutive epochs. All experiments are conducted over 10 random seeds, with the averaged results reported. Additional details can be found in Appendix A.3.
@@ 3.3.1. Ablation Studies
For the CoT reasoning, we randomly select user #3568 from Amazon-Book dataset, whose original profile is non-sensical, and we compare the new profiles generated with and without the application of four-step CoT reasoning, as shown in Fig. 7. It can be observed that when only given a prompt to regenerate the profile, the LLM merely attempts to modify the wording and structure of the original profile, yet still fails to deeply capture the user’s preferences. In contrast, the new profiles generated through four-step CoT reasoning not only greatly enrich the original descriptions but also introduce a substantial amount of new information (in purple color).
@@ 3.3.2. Hyper-parameter Sensitivities
Finally, we investigate the impact of two important hyperparameters, $|\mathcal{E}|$ and $|\mathcal{K}|$, as shown in Figs. 8 and 9, respectively. For the number of environments $|\mathcal{E}|$, LightGCN performs better with fewer environments (2–3), while Mult-VAE can accommodate a larger number of environments. For the number of profiles $|\mathcal{K}|$, both methods achieve the best performance with four or five profiles. An excessive number of profiles may not only introduce additional noise into the training process but also incur additional training costs.
@@ 4. Related Work
Figure 9. Impact of $|\mathcal{K}|$ for (a) LightGCN and (b) Mult-VAE w.r.t. Recall@20 and NDCG@20 on Amazon-Book dataset.
@@ Appendix A Appendix
In the Appendix, we first present the pseudo-code for the complete training process of the proposed ProEx. Subsequently, we provide a detailed introduction to all comparation methods and hyperparameter settings. Finally, we additionally provide experimental results to fully validate the effectiveness of ProEx.
1:  initialize model parameters for $f_{\Theta}$ and $f_{\psi}$;
3:  while ProEx not converge do
4:   sample a mini-batch of user-item pairs $\mathcal{B}$;
5:   for $u\in\mathcal{B}$ do
7:    for $k=1,2,...\mathcal{K}$ do
10:    end for
[…]
15:     construct fused representation ${\mathbf{z}}_{u}^{e}$ by Eq. 12;
16:     calculate the environment-specific loss $\mathcal{L}^{e}$ by Eq. 13;
17:    end for
19:   end for
20:   average gradients from mini-batch;
22:  end while
23:  return model parameters $\Theta,\psi$;
Algorithm 1 The training process of ProEx
A.1. Algorithm Description
The complete training procedure for the ProEx is described in Algorithm 1. Leveraging the LLM $f_{\text{LLM}}$ and text embedding model $f_{\text{text}}$, we first build multi-faceted profile $\mathcal{P}_{u}$ and $\mathcal{P}_{i}$ for each user $u$ and item $i$ that capture multiple aspects of user preferences (line 2). It is worth noting that this process is treated as a pre-processing step and is not involved in the training of the recommendation model.
In the training phase, a recommendation representation $\mathbf{z}_{u}$ is obtained for each user $u$ via the encoder of the recommendation model $f_{\Theta}$ (line 6). Subsequently, for each profile $s_{u,k}(k\in\mathcal{K})$, we generate an aligned representation $\tilde{\mathbf{c}}_{u,k}$ that matches the dimensionality of the recommendation representation $\mathbf{z}_{u}$ with the help of mapping function $f_{\psi}$ (lines 8-9). Recalling Eq. 6, we design two types of mapping functions: direct alignment $f_{\psi}^{d}$ and aggregation alignment $f_{\psi}^{a}$. For discriminative methods, we adopt the direct strategy $f_{\psi}^{d}$ to obtain representations for the user and item sides, respectively. Since many generative methods do not explicitly model item representations (liang2018variational; wang2023diffusion), we instead employ the aggregation strategy $f_{\psi}^{a}$ to construct an approximate distribution for the user $u$. To encourage diversity among the profiles, we apply a contrastive regularization loss $\mathcal{L}_{\text{reg}}$ on the aligned representations (line 11).
To fully leverage these profiles and mitigate potential biases, we introduce the concept of environments (line 12). Within each environment $e\in\mathcal{L}$, we assign a sampled coefficient $\vartheta_{k}^{e}$ to $k$-th profile (line 13) and then construct a mixed profile representation $\tilde{\mathbf{c}}_{u}^{e}$ through linear combination (line 14). Subsequently, a fusion function $f_{\gamma}$ is used to construct the final representation $\mathbf{z}_{u}^{e}$ for the recommendation task under the current environment $e$ (line 15). Finally, we compute the environment-specific loss $\mathcal{L}^{e}$ within each environment $e$, which consists of the recommendation loss $\mathcal{L}_{\text{rec}}^{e}$ and the alignment loss $\mathcal{L}_{\text{align}}^{e}$.
1:  load model parameters for $f_{\Theta}$ and $f_{\psi}$;
2:  for $u\in\mathcal{U}$ do
4:   for $k=1,2,...\mathcal{K}$ do
7:   end for
9:   construct fused representation ${\mathbf{z}}_{u}$ by Eq. 12;
11:  end for
12:  return Top-N recommendation list for each user $u\in\mathcal{U}$;
Algorithm 2 The inference process of ProEx
The complete inference procedure for the ProEx is described in Algorithm 2. The inference process follows a similar procedure to training. We first obtain the recommendation representation $\mathbf{z}_{u}$ of user $u$ using the recommendation model $f_{\Theta}$ (line 3). Subsequently, we perform the same mapping (line 5) and alignment process (line 6) for each profile. During the inference phase, we construct the mixed representation $\tilde{\mathbf{c}}_{u}$ using mean pooling instead of random weighting (line 8). Finally, we apply the fusion function $f_{\gamma}$ to construct the final representation $\mathbf{z}_{u}$ and predict the interaction probabilities between user $u$ and all items $\{i|i\in\mathcal{I}\}$ (line 10). It is worth noting that the inference phase does not incorporate multiple environments, and the item-side representations are modeled in a similar manner.
@@ A.1. Algorithm Description
A.2. Multi-faceted Profile Construction

### v2

@@ Abstract.
journalyear: 2026copyright: ccconference: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1; August 09–13, 2026; Jeju Island, Republic of Koreabooktitle: Proceedings of the 32nd ACM SIGKDD Conference on Knowledge Discovery and Data Mining V.1 (KDD ’26), August 09–13, 2026, Jeju Island, Republic of Koreadoi: 10.1145/3770854.3780284isbn: 979-8-4007-2258-5/2026/08ccs: Information systems Recommender systems
@@ 2.3.1. Multi-faceted Profile Construction
It is important to note that, since we aim to characterize the user from diverse perspectives and mitigate the representational collapse caused by popularity bias, we require the LLM in $\mathcal{F}_{4}$ to generate NPs that differ entirely from OP in terms of wording and expression. Upon completing the reasoning process, we can obtain the user $u$’s profile set $\mathcal{P}_{u}=\{s_{u,1},s_{u,2},...,s_{u,K}\}$, where $K=|\mathcal{P}_{u}|=|\mathcal{P}_{i}|$ is the number of profiles. $s_{u,1}$ denotes user $u$’s OP, and $\{s_{u,2},...,s_{u,K}\}$ correspond to the K1K-1 NPs of user $u$. For the item side, we have a similar definition for $\mathcal{P}_{i}=\{s_{i,1},s_{i,2},...,s_{i,K}\}$.
@@
(5) $\mathbf{C}_{u}=\{\mathbf{c}_{u,1},\mathbf{c}_{u,2},...,\mathbf{c}_{u,K}\},\ \text{where}\ \mathbf{c}_{u,k}=f_{\text{text}}(s_{u,k})\in\mathbb{R}^{1\times d_{s}}.$
@@ 2.3.2. Profile Extrapolation
In practice, we perform a quantitative analysis with $K=4$, where an original profile (i.e., OP) and three newly generated profiles (denoted as NP #1, NP #2, and NP #3) are obtained for each user based on multi-aspect profile construction. Fig. 4(a) illustrates the distribution of the original and newly generated profiles in the language space $\mathcal{Y}$. We observe that the original profiles (blue dots) are predominantly clustered around a central region (blue circle), reflecting a strong popularity bias. In contrast, the newly generated profiles are more widely dispersed throughout the feature space, minimizing their similarity to the original profiles. This observation demonstrates the effectiveness of the proposed CoT-based profile construction process. By occupying a broader region and expanding the feature space boundary, this process gives rise to profile extrapolation.
@@
(7) $\mathcal{L}_{\text{reg}}=\sum_{k=1}^{K}\text{log}\left(1+\text{exp}(1/\tau)\sum_{k^{\prime}=1,k^{\prime}\neq k}^{K}\text{exp}\left(\tilde{\mathbf{c}}_{u,k}^{\top}\tilde{\mathbf{c}}_{u,k^{\prime}}/\tau\right)\right),$
@@
Given the difficulty of evaluating the quality of profiles, it remains uncertain whether all newly generated profiles positively contribute to the recommendation. Based on this concern, we adopt Mult-VAE (liang2018variational) as the base model and incorporate each profile individually into its training process on Amazon-Book dataset (zhang2025towards), as shown in Fig. 4(b). As is evident, there exists a substantial performance disparity among different profiles, with NP #2 even leading to a significant degradation in recommendation performance. Combined with Fig. 4(a), although we explicitly enforce reduced semantic similarity in the newly generated profiles, this process may inadvertently compromise the original semantic information. Therefore, it is necessary to reconsider how to effectively integrate these profiles $\{\tilde{\mathbf{c}}_{u,k}\}_{k=1}^{L}$ for recommendation task.
@@
(8) $\tilde{\mathbf{c}}_{u}^{e}=\vartheta_{1}^{e}\tilde{\mathbf{c}}_{u,1}+\vartheta^{e}_{2}\tilde{\mathbf{c}}_{u,2}+...+\vartheta^{e}_{K}\tilde{\mathbf{c}}_{u,K};\quad\text{s.t.}\sum_{k=1}\vartheta_{k}^{e}=1,$
where LLM-generated profiles $\{\tilde{\mathbf{c}}_{u,k}\}_{k=1}^{K}$ are weighted by $\{\vartheta^{e}_{k}\}_{k=1}^{K}$.
For generative methods, according to Eq. 6, we model the mean $\tilde{\boldsymbol{\mu}}_{u,k}$ and variance $\tilde{\boldsymbol{\sigma}}^{2}_{u,k}$ over each user $u$’s $K$ profiles. Therefore, we consider them separately:
(9) $\tilde{\boldsymbol{\mu}}_{u}^{e}=\vartheta_{1}^{e}\tilde{\boldsymbol{\mu}}_{u,1}+\vartheta^{e}_{2}\tilde{\boldsymbol{\mu}}_{u,2}+...+\vartheta^{e}_{K}\tilde{\boldsymbol{\mu}}_{u,K};\quad\text{s.t.}\sum_{k=1}\vartheta_{k}^{e}=1,$
@@
In each environment $e$, the construction process for discriminative and generative methods is identical. Concretely, given a fixed set of profiles, their respective contributions are controlled by a set of coefficients $\{\vartheta^{e}_{k}\}_{k=1}^{K}$:
(10) $\{\vartheta_{1}^{e},\vartheta_{2}^{e},...,\vartheta_{K}^{e}\}\sim\text{Dirichlet}(\alpha_{1},\alpha_{2},...,\alpha_{K}),$
where $\{\alpha_{k}\}$ is a set of shape parameters (zhang2025mixrec) in the Dirichlet distribution. Here, we introduce a sampling-based strategy to generate entirely different combinations of coefficients $\{\vartheta^{e}_{k}\}_{k=1}^{K}$ in each environment $e$, thereby introducing variability across environments to enhance generalization.
@@
1:  initialize model parameters for $f_{\Theta}$ and $f_{\psi}$;
3:  while ProEx not converge do
4:   sample a mini-batch of user-item pairs $\mathcal{B}$;
5:   for $u\in\mathcal{B}$ do
7:    for $k=1,2,...\mathcal{K}$ do
10:    end for
[…]
15:     construct fused representation ${\mathbf{z}}_{u}^{e}$ by Eq. 12;
16:     calculate the environment-specific loss $\mathcal{L}^{e}$ by Eq. 13;
17:    end for
19:   end for
20:   average gradients from mini-batch;
22:  end while
23:  return model parameters $\Theta,\psi$;
Algorithm 1 The training process of ProEx
@@ 2.4. ProEx Framework
where $\lambda_{2}$ and $\lambda_{3}$ are also adjustable weights. For each profile $k\in\mathcal{K}$, we apply contrastive regularization $\mathcal{L}_{\text{reg}}$ (cf. Eq. 7) to further enable profile extrapolation, thereby exploring a broader preference space. Finally, we minimize the variance across all environments $\mathcal{E}$ to reduce inter-environment discrepancies, thereby extracting the invariant components of user preferences and maintaining a more stable training process (krueger2021out). The complete training and inference processes of ProEx are presented in Algorithm 1 and Algorithm 2, respectively.
1:  load model parameters for $f_{\Theta}$ and $f_{\psi}$;
2:  for $u\in\mathcal{U}$ do
4:   for $k=1,2,...\mathcal{K}$ do
7:   end for
9:   construct fused representation ${\mathbf{z}}_{u}$ by Eq. 12;
11:  end for
12:  return Top-N recommendation list for each user $u\in\mathcal{U}$;
Algorithm 2 The inference process of ProEx
Table 1. Statistics of the datasets.
Dataset #Users #Items #Interactions #Profiles
Amazon-Book 11,000 9,332 200,860 81,328
Yelp 11,091 11,010 277,535 88,404
Steam 23,310 5,237 525,922 114,188
@@ 3.1.1. Datasets
We utilize three widely adopted recommendation datasets — Amazon-Book, Yelp, and Steam (ren2024representation). Detailed statistics for three datasets are presented in Table 1. Each dataset is divided into training, validation, and test sets following a 3:1:1 split ratio. The last column shows the number of user and item profiles for each dataset, where each user (item) has one original profile (OP) and three new profiles (NP #1, NP #2, and NP #3).
@@ 3.1.4. Implementation Details
We implement ProEx by PyTorch [^2] https://github.com/BlueGhostYi/ProRec. All models are initialized using Xavier initialization (glorot2010understanding) and optimized with the Adam optimizer (kingma2014adam). The architectures of the base models adhere to the default configurations. The default batch size is 4,096 for discriminative methods, in contrast to 1,024 for generative methods. For all LLM-enhanced methods, we use OpenAI’s GPT-3.5-turbo as the base LLM and text-embedding-ada-002 (neelakantan2022text) for text embeddings to ensure a fair comparison. To evaluate the recommendation performance, we use the metrics Recall@N and NDCG@N (he2017neural). The full-ranking evaluation (he2020lightgcn) is applied for each user. Early stopping is activated when Recall@20 on the validation set does not improve for 20 consecutive epochs. All experiments are conducted over 10 random seeds, with the averaged results reported. Additional details can be found in Appendix A.2.3.
@@ 3.3.1. Ablation Studies
For the CoT reasoning, we randomly select user #3568 from Amazon-Book dataset, whose original profile is non-sensical, and we compare the new profiles generated with and without the application of four-step CoT reasoning, as shown in Fig. 7. It can be observed that when only given a prompt to regenerate the profile, the LLM merely attempts to modify the wording and structure of the original profile, yet still fails to capture the user’s preferences. In contrast, the new profiles generated via CoT reasoning not only greatly enrich the original descriptions but also introduce a substantial amount of new information (in purple color).
LightGCN Mult-VAE
R@20 N@20 R@20 N@20
base 0.1411 0.0856 0.1472 0.0916
3.3.2. Impact of Different Semantic Representations
Transforming the generated profiles into representations in the feature space is a crucial step in ProEx, which is primarily accomplished through the text embedding model $f_{\text{text}}$. We conduct experiments using various semantic representations on the base models LightGCN (he2020lightgcn) and Mult-VAE (liang2018variational), as shown in Table 5.
Firstly, all types of semantic representations improve the performance of the base models to varying degrees, further confirming the effectiveness and generalizability of ProEx. Secondly, although LLaMA3 produces the largest embeddings (4096), it achieves the worst performance, particularly with LightGCN. This may stem from the limited semantic understanding of the 8B-scale model, leading to low-quality representations. Moreover, the severe dimensional mismatch (e.g., 32 dimensions in LightGCN) makes alignment more difficult. Finally, the locally deployable SFR-Embedding-Mistral achieves strong performance, suggesting that ProEx does not necessarily rely on closed-source LLMs to be effective.
@@ 3.3.3. Hyper-parameter Sensitivities
In this section, we first investigate the impact of two important hyperparameters, $|\mathcal{E}|$ and $K$, as shown in Figs. 8 and 9, respectively. For the number of environments $|\mathcal{E}|$, LightGCN performs better with fewer environments (2–3), while Mult-VAE can accommodate a larger number of environments. For the number of profiles $K$, both methods achieve the best performance with four or five profiles. An excessive number of profiles may not only introduce additional noise into the training process but also incur additional training costs.
Finally, we investigate the weight of the contrastive regularization loss $\lambda_{2}$ used to enhance profile extrapolation, as shown in Fig. 10. It can be seen that Mult-VAE requires a larger loss weight, while LightGCN only needs a small range. This is expected: LightGCN uses a pairwise loss (rendle2009bpr) that is typically below 1, whereas Mult-VAE relies on a much larger reconstruction error (often tens to hundreds). Thus, a higher weight is needed to place the two losses on a comparable scale.
@@ 4. Related Work
Figure 9. Impact of $K$ for (a) LightGCN and (b) Mult-VAE w.r.t. Recall@20 and NDCG@20 on Amazon-Book dataset.
@@
(a) LightGCN
(b) Mult-VAE
Figure 10. Impact of $\lambda_{2}$ for (a) LightGCN and (b) Mult-VAE w.r.t. Recall@20 and NDCG@20 on Amazon-Book dataset.
@@ Appendix A Appendix
In the Appendix, we provide a detailed introduction to the prompt designs of multi-faceted profile construction, all comparation methods and hyperparameter settings.
@@
A.1. Multi-faceted Profile Construction